6x+6y=-6
5x+y=-13

Answer by oberobic(2304)

6x +6y = -6
5x + y = -13
.
divide the first equation by 6 and then subtract the second from it
.
x + y = -1
5x +y = -13
------------ subtract
-4x = 12
.
x = -3
.
substitute to find y
.
-3 +y = -1
y = 2
.
check these values using the second equation
5(-3) + 2 = -13
correct
.
Answer: x = -3, y = 2
.
Done.